# Whispering in the dark

### v2

@@ 5.2.1 HD96670
We assume the ratio of the terminal velocity to the escape velocity for O stars to be 2.6 (Vink et al. 2001a; Kudritzki & Puls 2000), =1 in the wind velocity law, and the Eddington factor at the surface of the O star to be 0.2 (see Eqs. 1-8 of Sen et al. 2021). Then, the estimated wind velocity of the O star at the position of the BH to be 932.3 km s-1. The orbital velocity is 375.4 km s-1. For the fiducial value of accretion efficiency (=1/3) and a non-spinning BH, we find that an accretion disc does not form (Eq. 10 of Sen et al. 2021). This is consistent with the lack of bright X-ray emission from this system.
We estimate the wind mass-loss rate to be one-third of the mass-loss rate derived from Eq. 24 of Vink et al. (2001b), 10-5.71 yr-1. The accretion radius, mass-accretion rate at the accretion radius , net mass-accretion rate near the BH event horizon and Eddington mass-accretion rate of the BH can be estimated to be 2.36 , 2.8510-9 yr-1, 4.8510-11 yr-1 and 1.3310-7 yr-1, respectively. For / = 0.00036, the radiative efficiency is 0.003 (figure 1 of Xie & Yuan 2012). The above quantities lead to an estimated X-ray luminosity of 81033 erg s-1 from ADAF. The observed X-ray flux from this system ranges from 2.21032 erg s-1 (NuSTAR, Gomez & Grindlay 2021) to 2.41034 erg s-1 (XMM-Newton, Saxton et al. 2008). Our results are in the right ballpark, although further study may be required to accurately constrain the X-ray variability and/or photometric properties of the system (e.g. see Wang & Grindlay 2022).
@@ 5.2.2 VFTS 243
The constrained stellar and binary parameters of VFTS 243 enable us to estimate the X-ray luminosity from ADAF around its 10 BH. The mass, radius and luminosity of the O star companion is 25 , 10.3 and respectively. The wind mass loss rate, with a clumping factor of 10, is 1.5 yr-1. For the adopted terminal wind velocity of 2100 km s-1, we estimate the net mass-accretion rate near the event horizon 1.6 yr-1. The Eddington mass-accretion rate of a non-spinning 10 BH is 2.2 yr-1, leading to a ratio 7.5. The radiative efficiency of an ADAF at the above ratio of is 0.0001, 0.001 and 0.01 (extrapolated from Fig. 1 of Xie & Yuan 2012) for = 0.001, 0.1 and 0.5 respectively. The resulting X-ray luminosity is 1029, and erg s-1 respectively. Hence, VFTS 243 is not expected to be observable even in faint X-rays. The Chandra T-ReX programme (Crowther et al. 2022) constrained the X-ray luminosity of this system log(/) ¡ 32.15, which is within our above estimates.
@@ 6.1 Disentangling thermal from non-thermal X-ray emission
Figure 5 shows the distribution of X-ray luminosity from ADAF and OB star luminosity for the BH+O star binaries that do not form a Keplerian accretion disc. For inefficient viscous heating (left panel), the peak of the distribution lies below the canonical wind X-ray emission = 10-7 relation (Crowther et al. 2022, solid black line). For efficient viscous heating (right panel), the peak of the distribution is nearly an order of magnitude above the typical wind X-ray luminosity expected from the main sequence companion to the BH. We suggest that targeted X-ray observations on the B stars in the SB1 sample of Shenar et al. (2022b) (see Table 1) provide a suitable opportunity to investigate the presence of BHs, and study accretion physics around stellar mass BHs. We identify two B stars, VFTS 186 and VFTS 640, in the T-ReX catalogue that show / ¿ 10-6. However, we note that Evans et al. (2015) categorises them as single stars based on radial velocity measurements. We identify two more O stars with / ¿ 10-6.5 in the T-ReX catalogue of Crowther et al. (2022, see their Table A2, T-ReX label cc4651, c7552) showing an X-ray excess, above the canonical wind X-ray luminosity. The X-ray excess could be ascribed to accretion onto an orbiting BH, a possibility which deserves follow-up observations. Finally, we notice that the non-detection of X-rays from the black hole candidate VFTS 243 sets an upper limit to the X-ray emission of , marginally above the expected emission from the wind (Shenar et al. 2022a).
Even in systems where the X-ray luminosity due to accretion is lower than the X-ray intrinsic emission from the stellar wind, detecting the former is not beyond reach. Indeed, their spectral properties are fundamentally different (e.g. Gierliński et al. 1999). In the radiatively inefficient regime, the X-ray luminosity from the low-density plasma near the BH is non-thermal (see Sect. 2.3.1). For instance, synchrotron emission from relativistic electrons accelerated by magnetic reconnection would yield a power-law of spectral index of 0 (El Mellah et al. 2022) to -0.7 (Ponti et al. 2017), depending on the efficiency of radiative cooling. In the case of a strong magnetic field, this emission could peak in hard X-rays, beyond the maxima from wind X-ray emission. This spectral disentangling between the thermal emission of the stellar wind and the non-thermal emission from the dilute accretion flow was used by Munar-Adrover et al. (2014) to measure an X-ray luminosity of from the X-ray faint system MWC 656, later refined to (Ribó et al. 2017). Even if the presence of the BH in this system has been challenged in a recent work (Janssens et al. 2023), it is a promising precedent which illustrates that non-thermal emission can be detected below the thermal wind X-ray luminosity level.
@@ 6.2 Be X-ray binaries and BH+OBe binaries
The typical X-ray luminosity of Be X-ray binaries is higher than that expected from advection-dominated accretion in the BH+OB star binaries. The predicted and the observed distribution of the outburst X-ray luminosity ranges from erg s-1 (Cheng et al. 2014 and Liu et al. 2024, Fig. 4 of). Hence, the X-ray luminosity of the majority of X-ray-faint BH+OB star binaries predicted in this work are not expected to overlap with the X-ray luminosity from the population of fairly brighter Be X-ray binaries or the BH+OB star binaries having a Keplerian accretion disc around the BH. We also note that all the known Be X-ray binaries are found to host a neutron star, not a BH (Ziolkowski 2014; Brown et al. 2018).
@@
Smoothed Particle Hydrodynamics simulations done by Brown et al. (2018) show that for an eccentricity of 0, the average X-ray luminosity due to mass accretion onto the BH from the decretion disc of the OBe star in a BH+OBe system is around erg (see their figures 6 and 7). We also note that Brown et al. (2018) assume a radiative efficiency of 0.1 in their X-ray luminosity estimates (see their Eq. 4), which may overestimate their X-ray luminosity prediction. Moreover, Zhang et al. (2004) propose that the decretion disc can get significantly truncated in the presence of a BH. Based on the above studies, the X-ray luminosity, due to mass accretion onto the BH from the decretion disc around the OBe star of a BH+OBe star binary, may not be strong enough to overlap with our predictions for X-ray luminosity from ADAF around the BH.
@@ 6.3 BH spin
Qin et al. (2018); Fuller & Ma (2019) showed that the black hole formed from the initially more massive star in a binary has a negligible spin parameter (see also Marchant et al. 2023). Moreover, a BH needs to accumulate mass equal to its mass to significantly increase its natal spin parameter (King & Kolb 1999; Wong et al. 2012), which is not feasible from sub-Eddington mass accretion (Fig. 1) over a timescale of a few Myrs. This is reflected in our assumption that the spin of the BH during the BH+OB star phase is zero. Sen et al. (2021) showed that the formation of an accretion disc is favoured in the case of more rapidly spinning black holes, as the innermost circular orbit radius of the black hole decreases with increasing BH spin. Hence, the predicted number of X-ray-bright BH+OB star binaries will increase for spinning black holes.
@@
Previous studies of X-ray-bright BH+OB star binaries such as Cyg X-1 (Gou et al. 2011; Miller-Jones et al. 2021; Zhao et al. 2021) and LMC X-1 (Gou et al. 2009; Mudambi et al. 2020) have inferred that the black holes in these systems are maximally spinning ( ¿ 0.90). On the other hand, recent simulations have found that deduced spins of wind-fed BH HMXBs may be model-dependent (Belczynski et al. 2021; Zdziarski et al. 2023). The X-ray spectral energy distribution can be statistically well-fitted with the assumption of a slowly spinning black hole and a Comptonised layer above the Keplerian accretion disc. Our population synthesis results, based on the accretion disc formation criteria of Sen et al. (2021), show that non-spinning BHs in BH+OB star binaries can form accretion discs around them in the LMC, relaxing the necessity for maximally spinning black holes to form wind-fed BH HMXBs. Furthermore, Batta et al. (2017) showed that a failed supernova explosion with sufficient fallback accretion can produce rapidly spinning BHs with a spin parameter ¿ 0.8, while the expected spin parameter from direct collapse is ¡ 0.3.
6.4 Accretion disc formation criterion
@@
Our criterion with overpredicts the number of observable wind-fed BH HMXBs in the LMC by an order of magnitude. For , Eq. (20) predicts 1.5% of the BH+OB binaries are expected to form an accretion disc. This leads to only 2 predicted X-ray-bright BH+OB star binaries in the LMC, consistent with one observed wind-fed BH HMXB, LMC X-1. Furthermore, for , the number of X-ray-faint BHs are 355180 for = 0.001, 0.1 & 0.5.
@@
Recent rapid binary population synthesis simulations by Romero-Shaw et al. (2023) showed that an accretion disc formation criterion that solely relies on the Roche lobe filling factor (Hirai & Mandel 2021) of the OB star companion may explain the scarcity of wind-fed BH HMXB in the Milky Way. However, their criterion was derived from a specific set of binary parameters, especially a mass ratio of 2 () and orbital periods up to d. At lower mass ratios, we find that the wind speed of an OB star companion to the BH may be low enough to form an accretion disc even at orbital periods of 10s of days (Fig. 3).
@@
If we additionally constrain our accretion disc formation criterion such that the OB companion must have a Roche lobe filling factor above 0.85 (see Fig. 3, we find that our models predict 2.5 X-ray-bright BH+OB star systems in the LMC. However, we do not see any physical reason to discount the BH+OB star binaries at long orbital periods that may also be observable as wind-fed BH HMXBs. We also note that the Roche lobe filling factor of the O star in HD96670 is 0.85, but no bright X-ray emission (1035 erg s-1) has been detected. A further investigation into the accretion disc formation may be useful, but beyond the scope of this work.
@@ 7 Conclusion
Understanding the production of X-rays from BH+OB star binaries is essential in order to explain the observed population of BH high-mass X-ray binaries (Fornasini et al. 2023; Zhang et al. 2024). In the present work, we study the population of BH+OB star binaries that are not expected to have a Keplerian accretion disc around the BH (Langer et al. 2020; Sen et al. 2021). These binaries are expected to be the higher mass counterparts of the population of Be X-ray binaries (Vinciguerra et al. 2020; Liu et al. 2024), but without copious X-ray emission (Sharma et al. 2007; Xie & Yuan 2012). Moreover, these X-ray-faint BH+OB star binaries are the tentative progenitors of the binary BHs that can merge within a Hubble time (Belczynski et al. 2008; Romero-Shaw et al. 2023) from the isolated binary evolution channel.
We use approximately 20 000 detailed binary evolution models that include differential rotation, time-dependent tidal interactions, and angular momentum transport during the mass-transfer phases. Only 8% of the total population of BH+OB star binaries in our grid are expected to form wind-fed BH HMXB systems like LMC X-1. Our models predict at least approximately 28 BH+OB star binaries to be observable in the LMC with current X-ray telescopes like Chandra and the upcoming SRG/eROSITA ( = - erg s-1).
We show in Sect. 2 that viscous heating imparted to electrons can be high in the presence of magnetic fields of 30 G, regardless of whether the flow is collisional (Xie & Yuan 2012) or not (Sharma et al. 2007). For efficient viscous heating, our models predict that a significant fraction of BH+OB star binaries without accretion discs can produce X-ray luminosity of - erg s-1 (Sect. 4.5). Up to about 72 X-ray-faint BH+OB star binaries in the LMC may be observable due to high viscous heating efficiency in ADAFs around stellar mass black holes.
Recent photometric and spectroscopic observations into the massive star content of the LMC provide an excellent testbed for our model predictions (Evans et al. 2011, 2015; Almeida et al. 2017; Mahy et al. 2020b, a). Identifying a population of X-ray-faint BH+OB star binaries from recent observational samples of SB1 binaries (Villaseñor et al. 2021; Shenar et al. 2022b; Mahy et al. 2022; Banyard et al. 2023) will provide crucial constraints on the contribution of the isolated binary evolution channel to the population of double BH mergers.
Our study of faint X-ray emission from BH+OB star binaries advocate for prolonged X-ray exposure of the stellar-mass BH candidates identified in the vicinity of OB stars without an X-ray-bright counterpart. Our method may also be applied to isolated BHs in order to produce complementary predictions in the X-ray band. This may be useful for future all-sky multi-wavelength surveys with high sensitivity to hunt for the large unidentified population of BHs in the Milky Way (Scarcella et al. 2021).

### v3

@@ 5.2.1 HD96670
We assume the ratio of the terminal velocity to the escape velocity for O stars to be 2.6 (Vink et al. 2001a; Kudritzki & Puls 2000), =1 in the wind velocity law, and the Eddington factor at the surface of the O star to be 0.2 (see equations 1-8 of Sen et al. 2021). Then, the estimated wind velocity of the O star at the position of the BH to be 932.3 km s-1. The orbital velocity is 375.4 km s-1. For the fiducial value of accretion efficiency (=1/3) and a non-spinning BH, we find that an accretion disk does not form (equation 10 of Sen et al. 2021). This is consistent with the lack of bright X-ray emission from this system.
We estimate the wind mass-loss rate to be one-third of the mass-loss rate derived from equation 24 of Vink et al. (2001b), 10-5.71 yr-1. The accretion radius, mass accretion rate at the accretion radius , net mass accretion rate near the BH event horizon and Eddington mass accretion rate of the BH can be estimated to be 2.36 , 2.8510-9 yr-1, 4.8510-11 yr-1 and 1.3310-7 yr-1, respectively. For / = 0.00036, the radiative efficiency is 0.003 (figure 1 of Xie & Yuan 2012). The above quantities lead to an estimated X-ray luminosity of 81033 erg s-1 from ADAF. The observed X-ray flux from this system ranges from 2.21032 erg s-1 (NuSTAR, Gomez & Grindlay 2021) to 2.41034 erg s-1 (XMM-Newton, Saxton et al. 2008). Our results are in the right ballpark, although further study may be required to accurately constrain the X-ray variability and/or photometric properties of the system (e.g., see Wang & Grindlay 2022).
@@ 5.2.2 VFTS 243
The constrained stellar and binary parameters of VFTS 243 enable us to estimate the X-ray luminosity from ADAF around its 10 BH. The mass, radius and luminosity of the O star companion is 25 , 10.3 and respectively. The wind mass loss rate, with a clumping factor of 10, is 1.5 yr-1. For the adopted terminal wind velocity of 2100 km s-1, we estimate the net mass accretion rate near the event horizon 1.6 yr-1. The Eddington mass accretion rate of a non-spinning 10 BH is 2.2 yr-1, leading to a ratio 7.5. The radiative efficiency of an ADAF at the above ratio of is 0.0001, 0.001 and 0.01 (extrapolated from Fig. 1 of Xie & Yuan 2012) for = 0.001, 0.1 and 0.5 respectively. The resulting X-ray luminosity is 1029, and erg s-1 respectively. Hence, VFTS 243 is not expected to be observable even in faint X-rays. The Chandra T-ReX programme (Crowther et al. 2022) constrained the X-ray luminosity of this system log(/) ¡ 32.15, which is within our above estimates.
@@ 6.1 Disentangling thermal from non-thermal X-ray emission
Figure 5 shows the distribution of X-ray luminosity from ADAF and OB star luminosity for the BH+O star binaries that do not form a Keplerian accretion disk. For inefficient viscous heating (left panel), the peak of the distribution lies below the canonical wind X-ray emission = 10-7 relation (Crowther et al. 2022, solid black line). For efficient viscous heating (right panel), the peak of the distribution is nearly an order of magnitude above the typical wind X-ray luminosity expected from the main sequence companion to the BH. We suggest that targeted X-ray observations on the B stars in the SB1 sample of Shenar et al. (2022b) (see Table 1) provide a suitable opportunity to investigate the presence of BHs, and study accretion physics around stellar mass BHs. We identify two B stars, VFTS 186 and VFTS 640, in the T-ReX catalogue that show / ¿ 10-6. However, we note that Evans et al. (2015) categorises them as single stars based on radial velocity measurements. We identify two more O stars with / ¿ 10-6.5 in the T-ReX catalogue of Crowther et al. (2022, see their Table A2, T-ReX label cc4651, c7552) showing an X-ray excess, above the canonical wind X-ray luminosity. The X-ray excess could be ascribed to accretion onto an orbiting BH, a possibility which deserves follow-up observations. Finally, we notice that the non-detection of X-rays from the black hole candidate VFTS 243 sets an upper limit to the X-ray emission of , marginally above the expected emission from the wind (Shenar et al. 2022a).
Even in systems where the X-ray luminosity due to accretion is lower than the X-ray intrinsic emission from the stellar wind, detecting the former is not beyond reach. Indeed, their spectral properties are fundamentally different (e.g., Gierliński et al. 1999). In the radiatively inefficient regime, the X-ray luminosity from the low-density plasma near the BH is non-thermal (see section 2.3.1). For instance, synchrotron emission from relativistic electrons accelerated by magnetic reconnection would yield a power-law of spectral index of 0 (El Mellah et al. 2022) to -0.7 (Ponti et al. 2017), depending on the efficiency of radiative cooling. In the case of a strong magnetic field, this emission could peak in hard X-rays, beyond the maxima from wind X-ray emission. This spectral disentangling between the thermal emission of the stellar wind and the non-thermal emission from the dilute accretion flow was used by Munar-Adrover et al. (2014) to measure an X-ray luminosity of from the X-ray faint system MWC 656, later refined to (Ribó et al. 2017). Even if the presence of the BH in this system has been challenged in a recent work (Janssens et al. 2023), it is a promising precedent which illustrates that non-thermal emission can be detected below the thermal wind X-ray luminosity level.
@@ 6.2 Be X-ray binaries and BH+OBe binaries
The typical X-ray luminosity of Be X-ray binaries is higher than that expected from advection-dominated accretion in the BH+OB star binaries. The predicted and the observed distribution of the outburst X-ray luminosity ranges from erg s-1 (Cheng et al. 2014 and Liu et al. 2024, Fig. 4 of). Hence, the X-ray luminosity of the majority of X-ray-faint BH+OB star binaries predicted in this work are not expected to overlap with the X-ray luminosity from the population of fairly brighter Be X-ray binaries or the BH+OB star binaries having a Keplerian accretion disk around the BH. We also note that all the known Be X-ray binaries are found to host a neutron star, not a BH (Ziolkowski 2014; Brown et al. 2018).
@@
Smoothed Particle Hydrodynamics simulations done by Brown et al. (2018) show that for an eccentricity of 0, the average X-ray luminosity due to mass accretion onto the BH from the decretion disk of the OBe star in a BH+OBe system is around erg (see their figures 6 and 7). We also note that Brown et al. (2018) assume a radiative efficiency of 0.1 in their X-ray luminosity estimates (see their equation 4), which may overestimate their X-ray luminosity prediction. Moreover, Zhang et al. (2004) propose that the decretion disk can get significantly truncated in the presence of a BH. Based on the above studies, the X-ray luminosity, due to mass accretion onto the BH from the decretion disk around the OBe star of a BH+OBe star binary, may not be strong enough to overlap with our predictions for X-ray luminosity from ADAF around the BH.
@@ 6.3 BH spin
Qin et al. (2018); Fuller & Ma (2019) showed that the black hole formed from the initially more massive star in a binary has a negligible spin parameter (see also Marchant et al. 2023). Moreover, a BH needs to accumulate mass equal to its mass to significantly increase its natal spin parameter (King & Kolb 1999; Wong et al. 2012), which is not feasible from sub-Eddington mass accretion (Fig. 1) over a timescale of a few Myrs. This is reflected in our assumption that the spin of the BH during the BH+OB star phase is zero. Sen et al. (2021) showed that the formation of an accretion disk is favoured in the case of more rapidly spinning black holes, as the innermost circular orbit radius of the black hole decreases with increasing BH spin. Hence, the predicted number of X-ray-bright BH+OB star binaries will increase for spinning black holes.
@@
Previous studies of X-ray-bright BH+OB star binaries such as Cyg X-1 (Gou et al. 2011; Miller-Jones et al. 2021; Zhao et al. 2021) and LMC X-1 (Gou et al. 2009; Mudambi et al. 2020) have inferred that the black holes in these systems are maximally spinning ( ¿ 0.90). On the other hand, recent simulations have found that deduced spins of wind-fed BH HMXBs may be model-dependent (Belczynski et al. 2021; Zdziarski et al. 2023). The X-ray spectral energy distribution can be statistically well-fitted with the assumption of a slowly spinning black hole and a Comptonised layer above the Keplerian accretion disk. Our population synthesis results, based on the accretion disk formation criteria of Sen et al. (2021), show that non-spinning BHs in BH+OB star binaries can form accretion disks around them in the LMC, relaxing the necessity for maximally spinning black holes to form wind-fed BH HMXBs. Furthermore, Batta et al. (2017) showed that a failed supernova explosion with sufficient fallback accretion can produce rapidly spinning BHs with a spin parameter ¿ 0.8, while the expected spin parameter from direct collapse is ¡ 0.3.
6.4 Accretion disk formation criterion
@@
Our criterion with overpredicts the number of observable wind-fed BH HMXBs in the LMC by an order of magnitude. For , Eq. (20) predicts 1.5% of the BH+OB binaries are expected to form an accretion disk. This leads to only 2 predicted X-ray-bright BH+OB star binaries in the LMC, consistent with one observed wind-fed BH HMXB, LMC X-1. Furthermore, for , the number of X-ray-faint BHs are 355180 for = 0.001, 0.1 & 0.5.
@@
Recent rapid binary population synthesis simulations by Romero-Shaw et al. (2023) showed that an accretion disk formation criterion that solely relies on the Roche lobe filling factor (Hirai & Mandel 2021) of the OB star companion may explain the scarcity of wind-fed BH HMXB in the Milky Way. However, their criterion was derived from a specific set of binary parameters, especially a mass ratio of 2 () and orbital periods up to d. At lower mass ratios, we find that the wind speed of an OB star companion to the BH may be low enough to form an accretion disk even at orbital periods of 10s of days (Fig. 3).
@@
If we additionally constrain our accretion disk formation criterion such that the OB companion must have a Roche lobe filling factor above 0.85 (see Fig. 3, we find that our models predict 2.5 X-ray-bright BH+OB star systems in the LMC. However, we do not see any physical reason to discount the BH+OB star binaries at long orbital periods that may also be observable as wind-fed BH HMXBs. We also note that the Roche lobe filling factor of the O star in HD96670 is 0.85, but no bright X-ray emission (1035 erg s-1) has been detected. A further investigation into the accretion disk formation may be useful, but beyond the scope of this work.
@@ 7 Conclusion
Understanding the production of X-rays from BH+OB star binaries is essential to explaining the observed population of BH high-mass X-ray binaries (Fornasini et al. 2023; Zhang et al. 2024). In this work, we study the population of BH+OB star binaries that are not expected to have a Keplerian accretion disk around the BH (Langer et al. 2020; Sen et al. 2021). These binaries are expected to be the higher mass counterparts of the population of Be X-ray binaries (Vinciguerra et al. 2020; Liu et al. 2024), but without copious X-ray emission (Sharma et al. 2007; Xie & Yuan 2012). Moreover, these X-ray-faint BH+OB star binaries are the tentative progenitors of the binary BHs that can merge within Hubble time (Belczynski et al. 2008; Romero-Shaw et al. 2023) from the isolated binary evolution channel.
We use 20 000 detailed binary evolution models that include differential rotation, time-dependent tidal interactions and angular momentum transport during the mass transfer phases. Only 8% of the total population of BH+OB star binaries in our grid are expected to form wind-fed BH HMXB systems like LMC X-1. Our models predict at least 28 BH+OB star binaries to be observable in the LMC with current X-ray telescopes like Chandra and the upcoming SRG/eROSITA ( = - erg s-1).
We show in Sect. 2 that viscous heating imparted to electrons can be high in the presence of magnetic fields of 30 G, whether the flow is collisional (Xie & Yuan 2012) or not (Sharma et al. 2007). For efficient viscous heating, our models predict that a significant fraction of BH+OB star binaries without accretion disks can produce X-ray luminosity between - erg s-1 (Sect. 4.5). Up to 72 X-ray-faint BH+OB star binaries in the LMC may be observable due to high viscous heating efficiency in ADAFs around stellar mass black holes.
Recent photometric and spectroscopic observations into the massive star content of the LMC provides an excellent testbed for our model predictions (Evans et al. 2011, 2015; Almeida et al. 2017; Mahy et al. 2020b, a). Identifying a population of X-ray-faint BH+OB star binaries from recent observational samples of SB1 binaries (Villaseñor et al. 2021; Shenar et al. 2022b; Mahy et al. 2022; Banyard et al. 2023) will provide crucial constraints on the contribution of the isolated binary evolution channel to the population of double BH mergers.
Our study of faint X-ray emission from BH+OB star binaries makes the case for long-time exposure surveys in the X-ray band on tentative BH+OB star candidate systems without an X-ray-bright counterpart. Our method may also be applied to isolated BHs to produce complementary predictions in the X-ray band. This may be useful for future all-sky multi-wavelength surveys with high sensitivity to hunt for the large unidentified population of BHs in the Milky Way (Scarcella et al. 2021).